# The Order Effect: Investigating Prompt Sensitivity in Closed-Source LLMs

### v1

@@ 4.1 Summary of Observations
  • Shuffling the input sequence almost always leads to accuracy degradation, likely due to the auto-regressive nature of LLMs. These models are trained to process inputs sequentially, so any disruption is treated as out-of-distribution that prevents performance gains.
  • For tasks involving well-defined and concise queries, LLMs tend to perform better. Queries such as multiple-choice in MMLU, are short, structured, and clearly communicate their objectives. Additionally, we assume that the presence of multiple options, rather than just two, reduces the randomness associated with binary decisions as observed in MRPC. For this type of dataset, LLMs seem to focus more on the task itself rather than the surface structure of the inputs. However, this remains a speculation based on the results we observed.
Regardless of the task or conditions, order sensitivity remains a significant challenge for LLMs. Despite ongoing advancements, these models continue to exhibit fragility when faced with something as simple as a shuffled input order. The fact that merely changing the order of choices in a question can consistently degrade performance, even in the strongest models, is a concerning limitation that demands further attention.
@@ 5 Conclusion and Future Work
In this paper, we investigated the problem of order sensitivity in LLMs, an issue that remains poorly understood despite prior studies. Our findings reveal a significant impact of input order on LLM performance, underscoring the need for further investigation. Understanding this sensitivity is crucial because LLMs are often used to generate metrics or act as automated judges in various pipelines, where inconsistent outputs can lead to misleading conclusions. Research relying on LLM-based evaluations may be affected by subtle factors such as input order which can impact outcomes.
@@
Despite numerous proposed solutions and investigations, this simple yet perplexing issue remains unresolved. Our future research will focus on addressing this problem, particularly within closed-source LLMs. In the near term, we plan to test non-GPT models on the same datasets and evaluate newer models with stronger reasoning capabilities, aiming to identify more definitive patterns. We hope that these additional data points will help streamline and enhance our investigation.
@@ Appendix A GPT-4o mini Responses
Figure 1: How GPT-4o mini responds to the same question when the order of choices is reversed. The calls were made on Wednesday, January 29th, at 11:54 a.m. EST.
@@ Appendix B Prompt Examples
In this example, the order of choices is kept untouched as this represents an original-order case. However, we added three examples to the prompt to ensure that our few-shot setup provides sufficient context for the users. Since there are three possible responses (A, B, or C), we provided one example for each option to help the system understand what “A” (the first option is preferred), “B” (the second option is preferred), and “C” (no preference) mean. Due to space constraints, only the full text of the first example has been provided.
@@
The prompt above includes a question with shuffled choices, meaning the first choice, “1) osteosarcoma”, is not necessarily the first choice in its original form within the dataset. Additionally, since this is a few-shot setup, we have included five examples alongside the original question to help communicate to the LLM what it is expected to do. The content of Examples 3 to 5 has been omitted for this example due to space constraints.

### v2

@@ 4.1 Summary of Findings
  • Shuffling the input sequence almost always led to decreased accuracy, likely due to the auto-regressive nature of LLMs. These models are trained to process inputs sequentially, so any disruption in that order is perceived as out-of-distribution, that prevents performance gains.
  • Few-shot learning was not as effective as anticipated.
The inability to identify a consistent pattern across tasks and settings highlights the severity and unpredictability of the order sensitivity problem. Regardless of task type or prompting strategy, input order remains an unresolved challenge for LLMs. The fact that merely reordering the choices in a question can consistently degrade performance, even in state-of-the-art models, is a serious limitation.
@@ 5 Conclusion and Future Work
In this paper, we investigated the problem of order sensitivity in LLMs, a phenomenon that remains poorly understood despite prior research. We included an analysis of the newly released DeepSeek model that is introduced well after the GPT series. We were expecting that it might exhibit different behavior, yet both older models (GPTs) and the newer one (DeepSeek) showed similar vulnerabilities to input order changes. Our findings highlight a significant impact of input order on LLM performance. This sensitivity is particularly important to understand, as LLMs are increasingly used to generate evaluation metrics or serve as automated judges in various pipelines. In these scenarios, inconsistent outputs can lead to misleading or unreliable conclusions. Research and applications that rely on LLM-based evaluations can be affected by subtle factors like input order, which can meaningfully influence results.
@@
Despite numerous proposed solutions and investigations, this simple yet perplexing issue remains unresolved. In the near term, we plan to test LLMs on a wider set of datasets and evaluate newer models with stronger reasoning capabilities. We also plan to investigate the impact of order in non-autoregressive LLMs to understand how much the architecture of current LLMs limits their ability to handle order-related issues.
@@ Appendix A GPT-4o mini Responses
Figure 1: How GPT-4o responds to the same question when the order of choices is reversed. The calls were made on Tuesday, May 6th, at 16:29 EST.
@@ Appendix B Prompt Examples
In this example, the order of choices is kept untouched as this represents an original-order case. However, we added three examples to the prompt to ensure that our few-shot setup provides sufficient context for the LLM. Since there are three possible responses (A, B, or C), we provided one example for each to help the system understand what “A” (the first option is preferred), “B” (the second option is preferred), and “C” (no preference) mean. Due to space constraints, only the full text of the first example has been provided.
@@
The prompt above includes a question with shuffled choices, meaning the first choice, “1) osteosarcoma”, is not necessarily the first choice in its original form within the dataset. Additionally, since this is a few-shot setup, we have included five examples alongside the original question to help communicate to the LLM what it is expected to do. The content of Examples 3 to 5 has been omitted due to space constraints.